# The trichotomy of primordial black holes initial conditions

### v1

@@ VI Initial conditions statistics
PBHs are formed by rarely large smoothed over-densities on a radius $r$. The threshold is then typically given in terms of the compaction function, which, if the perturbation is spherical, is the combination (4) of the statistical variable sheth
@@
where $r=|\vec{x}-\vec{x}_{0}|$ and $\zeta_{k}$ are assumed to be independent and Gaussian-distributed. Note that here we have not yet assumed any spherical symmetry and
@@
is the Fourier transformation of the Heaviside theta function. For spherically symmetric configurations of the over-density, as we already pointed out, $g(r,\vec{x_{0}})\propto\delta_{r}(\vec{x}_{0})$ and the combination (4) indeed corresponds to the effectively asymptotically flat “gravitational potential” of the perturbation.
Typically, one expects that a black hole is formed by a large over-density (peak) at the point $\vec{x}=\vec{x}_{0}$ which statistically implies an almost spherical symmetry around $\vec{x}_{0}$ bbks . In this case, without any loss of generality, one can fix $\vec{x}_{0}$ as the origin of the spherical symmetry.
@@
We then already see that in the case of Type-F configuration, where $\delta_{0}=0$, instead of a peak of the over-density we are in the presence of a trough (a minimum). Thus, statistically, Type-F configurations (or effectively Type-F) are hardly spherical michiro . In this case the threshold for PBH formation will be larger than the one discussed above, implying an extra statistical suppression for their realization with respect to the Type-C case. On the other hand, as we are going to see, the Type-F core is less statistically suppressed with respect to the Type-C/O ones.
@@
For illustrative purposes we will consider the two typical examples of a very narrow and a very broad $\mathcal{P}(k)$:
@@
In Fig.12 we see that at threshold, the core values for Type-C(I) are of $\sim\mathcal{O}(10^{-5})$ for all the cases studied.
Figure 12: Values of ${\cal R}(r)$ at the center for all the types of perturbations considered.
@@
To maximize $\sigma^{2}_{\delta_{0}}$ we consider the lowest asteroid mass PBH (unconstrained to be the whole of dark matter Carr ). In this case $M_{\bullet}\sim 10^{41}\text{ GeV}$ and so
@@
$\sqrt{\sigma_{\delta_{0}}^{2}}\ll\delta_{0}(\text{Type-C})\sim 10^{-5},$ (31)
implying a statistical suppression of those profiles with respect to Type-F.
For a very broad power spectrum however, the peak of the mass distribution might instead be related to the IR scale, as it would be in the presence of only Type-I black holes jacopo . In that case, for Type-C not to be statistically suppressed, one would need $k_{UV}/k_{IR}>10^{9}$. Then, the analysis of jacopo would follow since Type-C(I), given their lower threshold, would dominate the abundance of PBHs.
We note that in NANOGrav , such large separation of scales were considered when interpreting the NANOGrav signal nano as a second-order Gravitational Waves background sourced by large curvature perturbations.
@@ VII Conclusions
According to the separate universe approach, a local super-horizon patch (the core) of a perturbed Universe, is described by a FRW geometry with a certain three-dimensional curvature. At larger scales instead, when the inhomogeneities build up, the typical configuration related to a PBH formation is a shell with a over-density peak. The formation of a BH happens whenever the amplitude of this shell is high enough for the gravitational collapse to fight against the Universe dilution. If the core is a closed Universe (Type-C) configuration, it helps the collapse of the shell otherwise either it fights against it (Type-O(pen)) or it is irrelevant (Type-F(lat)). Therefore, for Type-F and O configuration, the required amplitude of the energy density to form a black hole (the threshold) is bound to be higher than that of Type-C. In this sense, there is a trichotomy of possible initial conditions, each one having a different threshold for BH formation depending upon the three-curvature peaks and cores amplitudes. In turn, this shows that the relevant variable for understanding BH formation in an expanding Universe is rather the non-linear over-density instead of the compaction function or the linear over-density, as thought earlier.
@@
This trichotomy is however only relevant if the core is extended enough to interfere with the shell. This empirically happens if the dimensionless curvature of the over-dense shell $w_{\cal R}\equiv-r^{2}\frac{\partial_{r}^{2}{\cal R}}{{\cal R}}$ calculated at the peak $r=r_{p}$, is higher than $\sim{\cal O}(17)$ ($\sim{\cal O}(30)$ for $w$), which is when the first point of Type-O crosses the Type-II region. For broader cases, the type of the black hole formed is always of Type-I. However, when the core becomes relevant, only Type-C configuration with large enough core amplitudes may generate PBHs of Type-I. In all other cases, Type-II black holes will always be formed. For Type-C, the separatrix between Type-I and Type-II core amplitudes has been empirically found in eq. (21), and depends upon the ratio of the core amplitude with respect to the peak one, being parameterized by $w_{\cal R}$. For very sharp shells, we found that the core amplitude must be at least about $70\%$ the peak one in order to form Type-I black holes.
@@
Finally, we argued that, at least for Gaussian distributed curvature perturbations, a zero curvature core (Type-F in spherical configuration) are more likely. Nevertheless, those cases are also luckily to generate a highly non-spherical initial condition for the non-linear over-density (note that this would be statistically different form the linear one). The reason is that a sharp shell can be thought as generated by the interference of an infinite number of peaks. In this case, the formation of a black hole would be harder than in the (more) spherical case of a Type-C with a sizable core. This happens because, for highly non-spherical configurations, large part of the energy density of the over-dense shell would be dissipated away in form of gravitational waves.
@@
For a very broad power spectrum however, which might be related to the NanoGrav gravitational waves observations NANOGrav , the statistical preference of Type-F with respect to Type-C becomes negligible. In this case then, the higher threshold necessary to form Type-F black holes would favor Type-C(I) for the PBHs abundance. This, in turn, would imply a black hole‘s mass spectrum peaked at the Infra-Red scale of the power spectrum jacopo , instead of at the Ultra-Violet scales, as typically assumed.
@@
Whether for a generic power spectrum the abundance of PBHs is finally dominated by Type-I black holes, as suggested in jacopo , or not, is however still an open question which, to be answered, requires further (physical and statistical) knowledge of the non-spherical gravitational collapse. This interesting problem is left for future studies.
@@ Appendix A Profile parameters
$w$ $w_{\cal R}$ $\lambda$ $\alpha$ $\beta$ $w$ $w_{\cal R}$ $\lambda$ $\alpha$ $\beta$
16.11 6.45 0.939 1 5.53 38.69 30.68 0.983 1.74 6
23.09 17.33 0.953 1 6.075 39.32 32.36 0.981 1.5 6.2
30.11 22.79 0.981 1.8 5.5 39.34 31.04 0.984 1.8 6
33.16 25.33 0.984 2 5.7 40.60 40.15 0.969 1 7.1
33.16 31.05 0.964 1 6.7 40.60 36.76 0.974 1.1542 6.7
33.81 23.42 0.998 13 5.1 42.10 37.74 0.976 1.2 6.7
34.08 24.88 0.990 3 5.4 42.32 34.28 0.983 1.6 6.2
34.12 24.44 0.993 4 5.3 42.71 38.10 0.976 1.2175 6.7
34.45 24.16 0.996 7 5.2 42.71 42.51 0.970 1 7.2
34.87 33.27 0.965 1 6.8 43.56 36.93 0.980 1.4 6.5
34.88 24.83 0.994 4.7 5.3 44.02 41.41 0.974 1.1 6.99
35.12 26.96 0.985 2 5.7 44.23 39.83 0.977 1.2 6.8
35.21 29.73 0.975 1.3 6.2 44.60 44.62 0.971 1 7.29
35.92 26.34 0.990 3 5.5 45.95 40.77 0.978 1.25 6.8
36.15 26.98 0.988 2.5 5.6 46.31 42.74 0.976 1.15 6.99
36.74 35.56 0.966 1 6.9 46.43 43.85 0.975 1.1 7.1
37.36 30.54 0.980 1.5 6.1 46.67 42.15 0.978 1.2 6.91
38.52 29.07 0.987 2.2 5.8 48.67 46.08 0.976 1.1 7.2
38.62 37.84 0.968 1 7
@@
Type $w$ $w_{\cal R}$ $\lambda$ $\alpha$ $\beta$
I 16.22 17.66 1.001 70 3.8
I 23.05 15.37 1.004 10 4.27
I 28.25 17.73 1.026 10 4.1
II 33.29 19.17 1.022 1 4.4
II 38.60 23.24 1.019 4.685
II 44.57 26.45 1.017 4.98
II 49.54 29.00 1.016 5.194
@@
$w$ $w_{\cal R}$ $\lambda$ $\alpha$ $\beta$ $w$ $w_{\cal R}$ $\lambda$ $\alpha$ $\beta$
34.02 23.51 0.998 15 5.1 45.63 37.97 0.983 1.5 6.5
34.74 24.29 0.996 7.8 5.2 46.21 34.16 0.992 3 6
35.25 25.10 0.994 5.1 5.3 46.29 40.95 0.979 1.26 6.8
35.96 25.85 0.993 4 5.4 47.55 41.58 0.980 1.3 6.8
36.86 26.81 0.992 3.4 5.5 47.96 40.86 0.982 1.4 6.7
37.08 26.91 0.992 3.5 5.5 47.98 43.62 0.978 1.19 6.99
37.88 27.85 0.991 3 5.6 48.39 43.83 0.978 1.2 6.99
41.17 31.97 0.987 2 6 48.48 35.82 0.993 3 6.1
42.34 33.97 0.984 1.7 6.2 48.51 35.00 0.995 4 6
43.48 34.53 0.985 1.8 6.2 48.56 42.71 0.978 1.2 6.9
44.25 37.28 0.981 1.43 6.5 48.63 44.05 0.978 1.2 7
44.65 37.49 0.982 1.45 6.5 48.80 45.14 0.977 1.15 7.1
45.30 39.54 0.979 1.3 6.7 49.68 46.61 0.977 1.12 7.2
45.44 35.42 0.988 2 6.2
Table 4: Parameters used for effective flat core profiles (Type-II) in Fig.7 and Fig.11 and the corresponding $w$ and $w_{\cal R}$.

### v2

@@ VI Initial conditions statistics
PBHs are formed by rarely large smoothed over-densities on a radius $r$ sheth . The threshold is then typically given in terms of the compaction function, which, if the perturbation is spherical, is the combination (4) of the statistical variable sheth
@@
where $r=|\vec{x}-\vec{x}_{0}|$ and $\zeta_{k}$, assumed to be independent and Gaussian-distributed, are the Fourier modes of the curvature perturbations in synchronous gauge and at leading order in gradient expansion. Note that here we have not yet assumed any spherical symmetry and
@@
is the Fourier transformation of the Heaviside theta function. For spherically symmetric configurations of the over-density, as we already pointed out, $g(r,\vec{x}_{0})=a_{i}H_{i}^{2}r^{2}\delta_{r}$, and the combination (4) indeed corresponds to the effectively asymptotically flat “gravitational potential” of the perturbation.
Typically, one expects that a black hole is formed by a large over-density (peak) at the point $\vec{x}=\vec{x}_{0}$, which statistically implies an almost spherical symmetry around $\vec{x}_{0}$ bbks . In this case, without any loss of generality, one can fix $\vec{x}_{0}$ as the origin of the spherical symmetry.
@@
We then already see that in the case of Type-F configuration, where $\delta_{0}=0$, instead of a peak of the over-density we are in the presence of a trough (a minimum). Thus, statistically, Type-C configurations (or effectively Type-F) are hardly spherical michiro . In this case, the threshold for PBH formation will be larger than the one discussed above, implying an extra statistical suppression for their realisation with respect to the Type-C case. On the other hand, as we are going to see, the Type-C(II) core is less statistically suppressed with respect to the Type-C(I) one.
@@
For illustrative purposes, we will consider the two typical examples of a very narrow and a very broad $\mathcal{P}(k)$:
@@
In Fig.10 we see that at threshold, the core values for Type-C(I) are of $\sim\mathcal{O}(10^{-5}-10^{-6})$ for all the cases studied.
Figure 10: Values of ${\cal R}(r)$ at the centre for Type-C perturbations.
@@
To maximise $\sigma^{2}_{\delta_{0}}$, we consider the lowest asteroid mass PBH (unconstrained to be the whole of dark matter Carr ). In this case $M_{\bullet}\sim 10^{41}\text{ GeV}$ and so
@@
$\sqrt{\sigma_{\delta_{0}}^{2}}\ll\delta_{0}(\text{Type-C(I)})\sim 10^{-5},$ (31)
implying a statistical suppression of Type-C(I) profiles with respect to Type-F, since one should go almost $10^{18}$ variances away from the central value to obtain Type-C(I) perturbations.
For a very broad power spectrum, however, the peak of the mass distribution might instead be related to the IR scale, as it would be in the presence of only Type-I black holes jacopo . In that case, for Type-C(I) not to be statistically suppressed, one would need $k_{UV}/k_{IR}>10^{9}$ to obtain $\sqrt{\sigma_{\delta_{0}}^{2}}\sim\delta_{0}(\text{Type-C(I)})$. Then, the analysis of jacopo would follow since Type-C(I), given their lower threshold, would dominate the abundance of PBHs.
We note that in NANOGrav , such a large separation of scales were considered when interpreting the NANOGrav signal nano as a second-order Gravitational Waves background sourced by large curvature perturbations.
@@ VII Conclusions
According to the separate Universe approach, a local super-horizon patch (the core) of a perturbed Universe is described by a FRW geometry with a certain three-dimensional curvature. At larger scales, instead when the inhomogeneities build up, the typical configuration related to a PBH formation is a shell with an over-density peak. The formation of a BH happens whenever the amplitude of this shell is high enough for the gravitational collapse to fight against the Universe dilution. If the core is a closed Universe (Type-C) configuration, it helps the collapse of the shell, otherwise, either it fights against it (Type-O(pen)), or it is irrelevant (Type-F(lat)). Therefore, for Type-F and O configurations, the required amplitude of the energy density to form a black hole (the threshold) is bound to be higher than that of Type-C. In this sense, there is a trichotomy of possible initial conditions, each one having a different threshold for BH formation depending upon the three-curvature peak and core amplitudes. In turn, this shows that the relevant variable for understanding BH formation in an expanding Universe is rather the non-linear over-density instead of the compaction function or the linear over-density, as thought earlier.
@@
This trichotomy is, however, only relevant if the core is extended enough to interfere with the shell. This empirically happens if the dimensionless curvature of the over-dense shell $w_{\cal R}\equiv-r^{2}\frac{\partial_{r}^{2}{\cal R}}{{\cal R}}$ calculated at the peak $r=r_{p}$, is higher than $\sim{\cal O}(17)$ ($\sim{\cal O}(30)$ for $w$), which is when the first point of Type-O crosses the Type-II region. For broader cases, the type of black hole formed is always of Type-I. However, when the core becomes relevant, only Type-C configuration with large enough core amplitudes may generate PBHs of Type-I. In all other cases, Type-II black holes will always be formed. For Type-C, the separatrix between Type-I and Type-II core amplitudes has been empirically found in eq. (21), and depends upon the ratio of the core amplitude with respect to the peak one, being parametrised by $w_{\cal R}$. For very sharp shells, we found that the core amplitude must be at least about $70\%$ the peak one in order to form Type-I black holes.
A more extensive and complete analysis should be done considering different perturbation profiles, $\zeta(r)$, which we leave for future work.
@@
Finally, we argued that, at least for Gaussian distributed curvature perturbations, shapes with zero curvature cores (Type-F in spherical configuration) are favourable. However, whether black holes will be more likely related to these shapes is a different question, which can only be answered with the statistics of thresholds. For example, Type-F configurations are also likely to generate a highly non-spherical initial condition for the non-linear over-density (note that this would be statistically different from the linear one), making gravitational collapse more difficult than in the spherical case. The reason is that a sharp shell can be thought of as generated by the interference of an infinite number of peaks, which would distort the spherical symmetry. For highly non-spherical configurations, a large part of the energy density of the over-dense shell would be dissipated away in the form of gravitational waves, thus a larger threshold.
@@
For a very broad power spectrum, however, which might be related to the NanoGrav gravitational waves observations NANOGrav , the statistical preference of Type-F with respect to Type-C(I) becomes negligible, in that case $\sqrt{\sigma_{\delta_{0}}^{2}}\sim\delta_{0}(\text{Type-C(I)})$. In this case, the higher threshold necessary to form Type-F black holes would favour Type-C(I) for the PBHs abundance. This, in turn, would imply a black hole’s mass spectrum peaked at the Infra-Red scale of the power spectrum jacopo , instead of at the Ultra-Violet scales, as typically assumed.
@@
Whether for a generic power spectrum, the abundance of PBHs is finally dominated by Type-I black holes, as suggested in jacopo , or not, is however still an open question which, to be answered, requires further (physical and statistical) knowledge of the non-spherical gravitational collapse. This interesting problem is left for future studies.
@@ Appendix A Profile parameters
$\mu$ $w$ $w_{\cal R}$ $\lambda$ $\alpha$ $\beta$ $\mu$ $w$ $w_{\cal R}$ $\lambda$ $\alpha$ $\beta$
0.678 16.11 6.45 0.939 1 5.53 0.605 38.69 30.68 0.983 1.74 6
0.671 23.09 17.33 0.953 1 6.075 0.616 39.32 32.36 0.981 1.5 6.2
0.613 30.11 22.79 0.981 1.8 5.5 0.602 39.34 31.04 0.984 1.8 6
0.600 33.16 25.33 0.984 2 5.7 0.658 40.60 40.15 0.969 1 7.1
0.665 33.16 31.05 0.964 1 6.7 0.634 40.60 36.76 0.974 1.1542 6.7
0.551 33.81 23.42 0.998 13 5.1 0.636 42.10 37.74 0.976 1.2 6.7
0.581 34.08 24.88 0.990 3 5.4 0.610 42.32 34.28 0.983 1.6 6.2
0.571 34.12 24.44 0.993 4 5.3 0.634 42.71 38.10 0.976 1.2175 6.7
0.558 34.45 24.16 0.996 7 5.2 0.658 42.71 42.51 0.970 1 7.2
0.662 34.87 33.27 0.965 1 6.8 0.619 43.56 36.93 0.980 1.4 6.5
0.565 34.88 24.83 0.994 4.7 5.3 0.645 44.02 41.41 0.974 1.1 6.99
0.600 35.30 27.02 0.985 2 5.7 0.635 44.23 39.83 0.977 1.2 6.8
0.633 35.21 29.73 0.975 1.3 6.2 0.657 44.60 44.62 0.971 1 7.29
0.578 35.92 26.34 0.990 3 5.5 0.629 45.95 40.77 0.978 1.25 6.8
0.586 36.15 26.98 0.988 2.5 5.6 0.638 46.31 42.74 0.976 1.15 6.99
0.661 36.74 35.56 0.966 1 6.9 0.644 46.43 43.85 0.975 1.1 7.1
0.618 37.36 30.54 0.980 1.5 6.1 0.634 46.67 42.15 0.978 1.2 6.91
0.590 38.52 29.07 0.987 2.2 5.8 0.643 48.67 46.08 0.976 1.1 7.2
0.659 38.62 37.84 0.968 1 7
@@
Type $\mu$ $w$ $w_{\cal R}$ $\lambda$ $\alpha$ $\beta$
I 0.595 16.22 17.66 1.001 70 3.8
I 0.558 23.05 15.37 1.004 10 4.27
I 0.457 28.25 17.73 1.026 10 4.1
II 0.444 33.29 19.17 1.022 1 4.4
II 0.433 38.60 23.24 1.019 4.685
II 0.422 44.57 26.45 1.017 4.98
II 0.417 49.54 29.00 1.016 5.194
@@
$\mu$ $w$ $w_{\cal R}$ $\lambda$ $\alpha$ $\beta$ $\mu$ $w$ $w_{\cal R}$ $\lambda$ $\alpha$ $\beta$
0.550 34.02 23.51 0.998 15 5.1 0.611 45.63 37.97 0.983 1.5 6.5
0.556 34.74 24.29 0.996 7.8 5.2 0.566 46.21 34.16 0.992 3 6
0.563 35.25 25.10 0.994 5.1 5.3 0.628 46.29 40.95 0.979 1.26 6.8
0.568 35.96 25.85 0.993 4 5.4 0.624 47.55 41.58 0.980 1.3 6.8
0.572 36.86 26.81 0.992 3.4 5.5 0.616 47.96 40.86 0.982 1.4 6.7
0.571 37.08 26.91 0.992 3.5 5.5 0.633 47.98 43.62 0.978 1.19 6.99
0.576 37.88 27.85 0.991 3 5.6 0.632 48.39 43.83 0.978 1.2 6.99
0.593 41.17 31.97 0.987 2 6 0.564 48.48 35.82 0.993 3 6.1
0.604 42.34 33.97 0.984 1.7 6.2 0.553 48.51 35.00 0.995 4 6
0.598 43.48 34.53 0.985 1.8 6.2 0.663 48.56 42.71 0.978 1.2 6.9
0.617 44.25 37.28 0.981 1.43 6.5 0.632 48.63 44.05 0.978 1.2 7
0.615 44.65 37.49 0.982 1.45 6.5 0.637 48.80 45.14 0.977 1.15 7.1
0.625 45.30 39.54 0.979 1.3 6.7 0.640 49.68 46.61 0.977 1.12 7.2
0.589 45.44 35.42 0.988 2 6.2
Table 4: Parameters used for Type-C(II) profiles in Fig.5 and Fig.9 and the corresponding $w$ and $w_{\cal R}$.
Appendix B Hamiltonian constraint
To test the convergence of the simulations, we have analysed the evolution of the Hamiltonian constraint equation, (as in albert )
$\mathcal{H}=D_{r}M-4\pi\Gamma\rho R^{2},$ (32)
[…]
The results for the evolution of the Hamiltonian constraint using the three profiles considered at threshold (Eq.(8)-(13)), are shown in Fig.11.
Figure 11: Numerical evolution of the Hamiltonian constraint for different profiles at threshold. The time of horizon crossing, when the perturbation re-enters the horizon, is $t_{H}=t_{0}\left(a_{i}r_{m}H_{i}\right)^{2}$, where $t_{0}=1$ for the simulations.
In Fig.11, one can see that around $t/t_{H}\sim 10^{1}$, the Hamiltonian constraint evolution oscillates and diverges from the previous monotonic behaviour. That is because around that point the BH is formed, and then the singularity, hence the numerical evolution does not hold any more.